### v1

# Selective Vision is the Challenge for Visual Reasoning: A Benchmark for Visual Argument Understanding

Abstract
Visual arguments, often used in advertising or social causes, rely on images to persuade viewers to do or believe something. Understanding these arguments requires selective vision: only specific visual stimuli within an image are relevant to the argument, and relevance can only be understood within the context of a broader argumentative structure. While visual arguments are readily appreciated by human audiences, we ask: are today’s AI capable of similar understanding?
We collect and release VisArgs, an annotated corpus designed to make explicit the (usually implicit) structures underlying visual arguments. VisArgs includes 1,611 images accompanied by three types of textual annotations: 5,112 visual premises (with region annotations), 5,574 commonsense premises, and reasoning trees connecting them to a broader argument. We propose three tasks over VisArgs to probe machine capacity for visual argument understanding: localization of premises, identification of premises, and deduction of conclusions. Experiments demonstrate that 1) machines cannot fully identify the relevant visual cues. The top-performing model, GPT-4-O, achieved an accuracy of only 78.5%, whereas humans reached 98.0%. All models showed a performance drop, with an average decrease in accuracy of 19.5%, when the comparison set was changed from objects outside the image to irrelevant objects within the image. Furthermore, 2) this limitation is the greatest factor impacting their performance in understanding visual arguments. Most models improved the most when given relevant visual premises as additional inputs, compared to other inputs, for deducing the conclusion of the visual argument. Our code and data are publicly released at https://github.com/JiwanChung/VisArgs.
Selective Vision is the Challenge for Visual Reasoning:
A Benchmark for Visual Argument Understanding
Jiwan Chungthanks:   denotes equal contribution  Sungjae Leefootnotemark:  Minseo Kim  Seungju Han♢♣ Ashkan YousefpourJack HesselYoungjae Yu $\spadesuit$ Yonsei University  $\diamondsuit$ Seoul National University  $\clubsuit$ Allen Institute for AI  $\heartsuit$ Samaya AI jiwan.chung@yonsei.ac.kr
1 Introduction
What we see depends
mainly on what we look for.
– Lubbock (1893)
Humans often communicate messages visually. For example, traffic light colors regulate drivers’ behavior, while computer icons, such as the trash bin symbol for deleting files or the magnifying glass for searching, guide user actions.
We consider the case of visual arguments. Consider Fig. 1, which depicts a polar bear on a shrinking ice floe. Without any text, this image calls attention to climate change: a visual metaphor connects melting ice to industrial emissions from factories. A plausible interpretation of the argument concludes: industrial pollution needs to be reduced.
Figure 1: An example from our VisArgs corpus. VisArgs makes the persuasion process in a visual argument explicit by representing it as a reasoning tree. Image credit: Eglė Plytnikaitė
Figure 2: To identify the bottleneck in visual argument understanding, we define three tasks over VisArgs: Localization of Premises requires models to ground the visual premises. Identification of Premises necessitates models to infer the visual premise relevant to the given intermediate conclusion. Deduction of Conclusion studies the ability of models to deduce the argument’s conclusion based on different levels of inputs.
We introduce VisArgs, an annotated dataset of 1,611 images containing visual arguments. VisArgs makes explicit the reasoning process in interpreting a visual argument: [^1] We note that our corpus contains just one possible interpretation of a visual argument (rather than, e.g., claiming to represent the creator’s intent). each image is annotated with visual premises grounded on object bounding boxes, commonsense premises eliciting implicit knowledge, and argument trees formalizing the connection of these premises to the conclusion. An argument tree consists of a root node (conclusion), some internal nodes (intermediate conclusion), and two types of leaf nodes (visual and commonsense premises).
Using VisArgs, we propose three complementary tasks to evaluate different aspects of machine capacity for comprehending visual arguments as illustrated in Fig. 2: 1) Localization of Premises: associates the description of a visual premise with a specific region in the image, 2) Identification of Premises: Given an image and an (intermediate) conclusion, retrieves the necessary visual premises to support the conclusion, and 3) Deduction of Conclusion: generates the conclusion with increasing detail of the annotated visual argument.
Experiments on VisArgs demonstrate that the main bottleneck for machine understanding of visual arguments is selective vision, i.e., Identification of Premises relevant to a given conclusion (see § 5.2). We show that while machines can identify visual premises within an image (albeit worse than human agreement, see Localization of premises § 5.1), they struggle to discern which premises are relevant to the conclusion among them. Results on our final Deduction of Conclusion task (§ 5.3) additionally support the hypothesis that difficulties in understanding visual arguments do not stem from deficiencies in raw vision capacity. There, we controlled the level of input to the algorithm, ranging from raw images to explicit reasoning trees. The greatest accuracy gains came from the inclusion of relevant visual cues, further supporting our main hypothesis. In all visual argument understanding tasks, machines perform worse than human agreement, providing avenues for future work.
In conclusion, our results suggest that selective attention to visual cues is the main bottleneck for the current AI capacity to understand visual arguments. This finding also establishes visual argument understanding as a distinct area of study in the computational domain: vision does not precede, but works jointly with reasoning in terms of understanding visual arguments. We expect that VisArgs will be utilized as a diagnostic benchmark for selective vision in future multimodal models: even the best current models lag significantly behind human performance in our Identification of Premises and Deduction of Conclusion tasks.
2 Related Work
Visual arguments are arguments built on visual medium Boland (2005). Unlike typical images, a visual argument is intentionally organized to persuade viewers to a certain conclusion Birdsell and Groarke (1996); Boland (2005). This work builds upon to ongoing debates in the human studies literature about the nature of visual arguments Johnson (2003); Tseronis (2018). Our results (§ 5) suggest that understanding visual arguments requires focusing on a subset of the visual context: not all visual cues contribute, and identifying the relevant ones is the key necessity. This task one of selective vision: the human capability to focus on behaviorally relevant stimuli. Desimone and Duncan (1995). Examples of visual arguments are prevalent in advertisements Kjeldsen (2012); Zhang et al. (2018); Ye et al. (2019), cartoons Birdsell and Groarke (2007), mathematical educations Inglis and Mejía-Ramos (2009), and, arguably, diagrams Kembhavi et al. (2016); Alikhani and Stone (2018).
Multimodal reasoning. Recent studies have introduced various multimodal models capable of sophisticated reasoning across different modalities, such as vision and language. Models such as LLaVA Liu et al. (2023a), Idefics2 Laurençon et al. (2024), and Qwen-VL Bai et al. (2023) are built on pretrained large language models (e.g., LLaMA Touvron et al. (2023)) and integrate vision encoders. Others, including OFA Wang et al. (2022) and Unified-IO Lu et al. (2022), are developed from scratch. These models excel in tasks such as localization, image captioning, and commonsense reasoning. Furthermore, models such as Unified-IO-2 Lu et al. (2023) and GPT-4-O Achiam et al. (2023) can understand audio, while others Zellers et al. (2022); Han et al. (2023a) support video understanding, demonstrating broad multimodal reasoning capabilities.
Beyond factual visual understanding. Visual comprehension is moving beyond factual understanding to include various types of writing. These include visual commonsense reasoning Zellers et al. (2019); Park et al. (2020); Han et al. (2023b); Hessel et al. (2022), humor understanding Hessel et al. (2023); Hyun et al. (2023), and understanding social interaction Zadeh et al. (2018). Of particular relevance to our work is visual metaphors Akula et al. (2023), which express abstract concepts with concrete visual cues. While some overlap exists in the images used, there are clear differences in intention and structure; not all metaphorical images present clear arguments and can be seen as visual arguments. Conversely, not all visual arguments depend on metaphors Blair (2012).
Argument structure. An argument is typically understood as a structure that starts from a set of premises (reasons) and ends in a conclusion, often represented symbolically as a tree Whately (1863); Freeman (2011). While there have been extensions, including computational models of arguments Bench-Capon and Dunne (2007); Rahwan and Simari (2009); Atkinson et al. (2017), we use the basic form of trees connecting premises to conclusions, following previous literature Stab and Gurevych (2014); Lawrence and Reed (2020).
3 VisArgs Dataset
Figure 3: Human workers iteratively refine initial data produced by machines in VisArgs annotation process.
VisArgs comprises a total of 1,611 images featuring clear visual arguments. These images are categorized into 914 advertisement images and 697 cartoon images based on their sources. Each image in VisArgs is annotated with descriptions and bounding boxes for the visual premises (VP), descriptions of the commonsense premises (CP), the conclusion, and an argumentation tree (T) detailing the reasoning path from the premises to the conclusion (C). All descriptions are in English, with an average character length of 79, 91, 142, and 105 for VP, CP, C, and T, respectively. On average, each image contains 3.17 visual premises, 3.46 commonsense premises, and 2.88 intermediate conclusions.
3.1 Annotation Process
We partially rely on GPT-4-O Achiam et al. (2023) for initial annotations. However, these machine-generated annotations serve only as preliminary seeds, which are then extensively refined by experienced human workers, as illustrated in §3. The machine’s role is merely to provide imperfect starting points to facilitate the human annotation process. Below, we detail our annotation procedure.
Collecting Images. We manually collect around 1,600 images from Pinterest.222www.pinterest.com Starting with keyword-based searches (e.g. creative ads), we expanded our collection by exploring related images. Cartoons (which often contain visual arguments Birdsell and Groarke (1996)) were sourced from a dedicated website.333www.cartoonmovement.com We manually collected around 1,600 cartoons from various categories, including politics, education, and environment. For both categories, we followed previous work Schuhmann et al. (2022); Lee et al. (2021) by including URLs to the images to comply with licensing terms. Refer to Appendix A for details.
Describing Visual Premises. The next step is to explicitly describe the visual argument within each image. However, during the early stages of our annotation process, we discovered that although humans can naturally understand visual arguments, they often find it challenging to articulate their interpretation into structured argumentation trees. Therefore, we used an AI model (GPT-4-O) to generate initial candidates. Human workers then select and modify these initial annotations, as shown in Fig. 3. To facilitate this process, we break down the annotation into two steps: describing the visual premises and specifying the argument structure.
Given an image containing a visual argument, we instructed the model to generate a set of visual premises necessary to support the argument (refer to Appendix J for further details). However, the AI model often fails to fully comprehend the visual argument. To address this, we engaged a pool of experienced human workers to review the machine-generated outputs. They selected the correct visual premises and made necessary modifications to ensure accuracy and coherence. Additionally, we identified that a model-generated visual premise sometimes contains multiple atomic premises. We instructed the reviewers to separate these merged premises into individual atomic premises. Further details are provided in Appendix A.
Specifying Argument Structure. Given the visual premises and the image, we further annotate three components constituting the argumentation structure: commonsense premises, conclusions, and argument trees. As in the previous stage, we first generate initial candidates using an AI model. For this stage, we impose an additional criterion: the set of selected premises should be both necessary and complete (refer to Appendix J). The same pool of human workers then adjust the annotations for greater accuracy. The workers first verify the correctness of the conclusion and discard the image if it is incorrect. They then identify and correct any errors, including semantic and structural mistakes. We discarded 1,593 of the 3,204 images in this process. Details are provided in Appendix A.
Visual Grounding. Lastly, we manually gather bounding box annotations for each visual premise to finalize the multimodal annotations. We assume a one-to-one relationship between each bounding box ($vp^{r}_{i}$) and its corresponding textual description ($vp^{d}_{i}$). Annotators are instructed to ensure accurate matching and precise bounding box tightness, as detailed in Appendix A.
3.2 Data Analysis
Topic Diversity. To gauge the diversity of topics covered in VisArgs, we run zero-shot categorization using GPT-4-O and LLaMa3 AI@Meta (2024) to classify the topics of visual premises and conclusions. The topics cover a wide range of visual objects and argument topics, as shown in Fig. 4. Refer to Appendix B for details.
Figure 4: Variety of the topics represented in the visual premises and conclusions in VisArgs.
Recall Hit rate
LLaVaNeXT 0.48 0.14
LLaVa-LLaMa3-Docci 0.27 0.02
ShareCaptioner 0.40 0.12
Table 1: Frequency of detailed captions containing visual premises. Hit rate denotes how often all visual premises per image are included in the captions.
Visual Cues vs. Dense Captioning. In theory, selective attention to visual premises could be collapsed into an NLP problem by describing everything in an image. To test this counter-hypothesis, we manually check how often the visual premises are contained in the outputs of detailed captioning models. We include three baselines here: a generalist (LLaVA-Next Liu et al. (2024b)), a specialist (ShareCaptioner Chen et al. (2023)), and LLaVA-LLaMa3 XTuner Contributors (2023) fine-tuned on a detailed captioning corpus (DOCCI Onoe et al. (2024))444huggingface.co/gokaygokay/llava-llama3-docci. Tab. 1 summarizes our manual inspection of 100 images, showing that the detailed captions insufficiently capture the visual premises, with the hit rate staying below $15\%$ for all models.
Safety. Since we did not initially filter for safety, we now analyze the safety of VisArgs using standard models. For textual safety, we utilize the Perspective API555www.perspectiveapi.com; June 2024 version., and for visual domains, we employ LAION-Safety666www.github.com/LAION-AI/LAION-SAFETY. The toxicity scores for textual descriptions were $0.03$ for visual premises and $0.07$ for conclusions. Also, given the threshold of $0.7$, no descriptions and visual premises were classified as toxic. Furthermore, only $71$ among $1611$ images are classified as unsafe. Manual inspection reveals that such “unsafe" images were social campaigns advocating against the harmful behaviors which presumably triggered the LAION detector.
4 Task Overview
We pose three tasks based on VisArgs for a structured analysis of how machines understand arguments presented in visual form.
Acc. Prec. Rec. F1 Corr. ($\rho$)
BLEU-4 67 44 67 53 18
ROUGE 75 76 75 72 35
CIDEr 72 70 72 70 26
GPTEval 75 83 75 76 53
BERTScore 94 94 93 93 59
Table 2: Correlation of each metric with human decisions in the Deduction of Conclusion task.
[…]
4.1 Localization of Premises
The first task focuses on assessing whether machines can accurately align visual premises ($\textsc{VP}^{d}$) with the corresponding regions ($\textsc{VP}^{r}$) in a given image ($I$), requiring minimal computational reasoning capabilities. It aims to determine if difficulties in understanding visual arguments originate from basic object detection stages.
We investigate two setups based on the algorithm’s ability to output bounding box labels: First, closed-set grounding is designed for a broad range of models that lack explicit grounding capabilities. The problem is formulated as a retrieval task where the goal is to match a region in the image ($vp^{r}_{i}$) with an appropriate description ($vp^{d}_{i}$). We adapt standard image-text matching models (e.g. CLIP) to perform grounded image-text matching. More details can be found in § 5. Second, open-set grounding tests models with explicit grounding capabilities. The task is framed as a visual grounding problem Yu et al. (2016), where the machine must locate an object in an image based on a natural language expression. Both the ground truth and machine output are represented as bounding box coordinates $\langle x,y,h,w\rangle$. Performance is evaluated using the intersection over union (IoU) ratio, with predictions considered correct if IoU $\geq 0.5$.
4.2 Identification of Premises
The second task tests the machines’ capabilities to discern visual premises that would better support the given conclusion. Given the image $I$, the intermediate conclusion $ic$, and a superset of the gold text descriptions of the visual premises $S\supset\textsc{VP}^{d}$, the machine should retrieve a correct visual premise $vp^{d}_{i}\in\textsc{VP}^{d}$. Note that the candidate set $S$ contains a single ground truth premise $vp^{d}_{i}$ and a fixed number $K=2$ of negative premises.
The complexity of a retrieval task is impacted by the choice of the negative set. We explore four types of global samplers and a single local sampler for constructing the negative set. The global samplers source the negatives from visual premises that do not correspond to the selected image. The only difference is the sample selection strategy: 1. Random sampling samples uniformly without replacement. 2. Visual sampling samples from the top premise descriptions that are the closest to the given image. We use CLIPScore Hessel et al. (2021) for the multimodal scoring. 3. Textual sampling samples from the top premise descriptions that are the closest to the ground truth premise. We use cosine similarity on the ColBERT Khattab and Zaharia (2020) representation space for the textual scoring. 4. Mixed sampling combines textual and visual sampling by visually selecting from the top 10 textual retrieval results.
For local sampling, we select from the visual premises that do correspond to the given image. Relying on our argumentation tree annotation, we can automatically obtain the set of local visual premises that does not help justify the given intermediate conclusion $ic$. we sample uniformly without duplicates from the local pool and name the method 5. Semantic sampling due to its argumentation-dependent nature. Additionally, we report human performance on 100 random samples to mitigate the risk of false negatives.
4.3 Deduction of Conclusion
The final task is to evaluate how each component ($I$, VP, CP, IC, and $T$) influences the deduction of the conclusion C. We approach this as a sequence-to-sequence task aimed at generating C. While this allows flexible output formats, it complicates evaluation because the machine-generated text must be compared to the free-form label. Common text comparison practices, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and CIDER Vedantam et al. (2015) measure surface form similarity, not semantic similarity between conclusions. Alternatively, prompt-based evaluation using general reasoners (e.g. GPT-4) Achiam et al. (2023) can be biased by factors including candidate order Pezeshkpour and Hruschka (2023). Human verification, though ideal, is costly and hard to reproduce. We conduct a small-scale comparison study (see Tab. 2) to verify that the model-based metric BERTScore Zhang* et al. (2020) provides the most stable estimate, making it our primary metric. Details are in Appendix D.
[…]
Ads Cartoon All
Random 33.33 33.33 33.33
Human 100.00 100.00 100.00
CLIPRN50 80.83 82.72 81.91
CLIPViT-L 82.72 82.96 82.85
CLIPViT-L@336 82.09 83.26 82.76
SigLIP 86.10 86.67 86.43
AlphaCLIP 75.15 77.44 76.45
OFABase 68.75 75.71 72.71
OFALarge 72.01 79.18 76.10
Table 3: closed-set results in localization of premises.
[…]
UNINEXT-H 38.75 35.58
LISA 44.25 44.62
Unified-IO-2 48.61 47.15
OFA 50.14 49.13
MM-G-Dino 55.02 54.98
Table 4: open-set results in localization of premises.
Global Local
Random Visual Textual Mixed Semantic
Random 33.33 33.33 33.33 33.33 33.33 (-)
Human 100.00 99.00 94.00 100.00 98.00 ($\uparrow$ 4.00)
OFA 0.00 0.00 0.00 0.00 0.00 (-)
Qwen-VL-Chat 86.05 85.77 70.67 75.57 49.74 ($\downarrow$ 20.93)
CogVLM 97.46 96.39 88.00 92.22 65.31 ($\downarrow$ 22.69)
Idefics2 98.68 97.83 91.80 95.07 75.01 ($\downarrow$ 16.79)
InstructBLIP 83.77 79.23 66.95 71.37 61.90 ($\downarrow$ 5.05)
Unified-IO-2 98.42 96.99 86.87 92.81 34.74 ($\downarrow$ 52.13)
LLaVA-1.5 98.65 97.91 83.74 89.86 67.43 ($\downarrow$ 16.31)
LLaVA-NeXT 97.66 96.20 80.90 85.86 78.53 ($\downarrow$ 2.37)
GPT-4-O - - - - 79.50 (-)
Semantic
+ G.T region
78.13 ($\uparrow$ 16.23)
84.39 ($\uparrow$ 49.65)
76.67 ($\uparrow$ 9.24)
82.19 ($\uparrow$ 3.66)
Table 5: Results of the Identification of Premises task. Difference between the lowest score in global and local setup for each model are highlighted.
5 Experiments
5.1 Localization of Premises
Localization of Premises tests the visual grounding capabilities of machines. Given the image $I$ and description of a visual premise $vp^{d}$, the goal is to find a corresponding region $vp^{r}$ in the image.
Metrics and Models. For closed-set grounding, which is an N-way classification task, the goal is to match the given description with the correct bounding box. To evaluate standard image-text matching algorithms (e.g. CLIP), we crop the regions accordingly. The models for this task include various CLIP-based models (CLIP Radford et al. (2021) with different backbones and SigLIP Zhai et al. (2023)) and a multitask model OFA Wang et al. (2022). For open-set grounding, which is to locate an object in an image based on a natural language expression, we instruct the models to output bounding box coordinates and we compare them to the ground truth region. A predicted coordinate is considered correct if its intersection over union with the gold label is at least (IoU $\geq 0.5$). We use a diverse set of models that support local region output formats, UNINEXT-H Yan et al. (2023), LISA Lai et al. (2023), Unified-IO-2 Lu et al. (2023), OFA, MM-G-DINO Liu et al. (2023b).
Results. Tab. 3 demonstrates that current models are generally effective in matching descriptions of visual premises to the correct regions in images, thereby meeting the basic vision requirements for understanding visual arguments. However, the results for open-set grounding, shown in Tab. 4, are somewhat mixed: the scores are acceptable but not uniformly high. We traced this performance decline to the nature of zero-shot object detectors, which are designed to detect concrete objects and clear segments. In contrast, our bounding boxes are more semantic Guo et al. (2018). Visual examples can be found in Appendix G.
Image + VP + CP + Tree
LLaMA3 - 30.2 37.8 ($\uparrow$7.6) 40.8 ($\uparrow$2.0)
Mistralv0.2 - 18.9 30.2 ($\uparrow$11.3) 36.6 ($\uparrow$6.4)
Zephyr - 20.6 28.7 ($\uparrow$8.1) 36.5 ($\uparrow$7.8)
OFA -41.3 -24.6 ($\uparrow$16.7) -16.5 ($\uparrow$8.1) -13.9 ($\uparrow$2.6)
Qwen-VL-Chat 12.8 23.7 ($\uparrow$10.9) 30.2 ($\uparrow$6.5) 32.7 ($\uparrow$2.5)
CogVLM 25.7 30.7 ($\uparrow$5.0) 33.6 ($\uparrow$2.9) 36.3 ($\uparrow$2.7)
Idefics2 16.4 22.8 ($\uparrow$6.4) 29.5 ($\uparrow$6.7) 36.6 ($\uparrow$7.2)
InstructBLIP -18.4 16.6 ($\uparrow$35.0) 28.9 ($\uparrow$12.3) 32.2 ($\uparrow$3.3)
Unified-IO-2 -9.9 -3.4 ($\uparrow$6.5) 4.2 ($\uparrow$7.6) 8.0 ($\uparrow$3.8)
LLaVA-1.5 2.2 20.0 ($\uparrow$17.8) 29.6 ($\uparrow$9.6) 33.7 ($\uparrow$4.1)
LLaVA-Next 15.1 28.4 ($\uparrow$13.3) 34.3 ($\uparrow$5.9) 39.5 ($\uparrow$5.2)
GPT-4-O 25.5 - 34.3 ($\uparrow$8.8) 41.0 ($\uparrow$6.7)
Table 6: Results of the Deduction of Conclusion task, showing how incremental additions of inputs affect the correctness of the conclusion. Scores are presented using BERTScore, with similar trends observed across other metrics as detailed in Appendix F.
5.2 Identification of Premises
Identification of Premises tests the selective attention capabilities, i.e., selecting necessary visual cues to understand an argument. Given the image $I$ and an intermediate conclusion $ic$, the goal is to select a visual premise $vp^{d}$ that leads to this intermediate conclusion.
Metrics and Models. For this task, we retain only intermediate conclusions that have at least two unrelated visual premises within the image. We report classification accuracy based on a single gold visual premise and two negative candidates. The negative sets are sourced as described in § 4.2 and are categorized into random, visual, textual, mixed, and semantic sets. Given the task’s requirement for understanding argumentation structure, the models evaluated are primarily multimodal large language models with adequate reasoning capabilities. We experiment with a broad selection of models: OFA Wang et al. (2022), Qwen-VL-Chat Bai et al. (2023), CogVLM Wang et al. (2023), Idefics2 Laurençon et al. (2024), InstructBLIP Dai et al. (2024), Unified-IO 2 Lu et al. (2023), LLaVa-1.5 Liu et al. (2024a), and LLaVa-Next Liu et al. (2024b). For the sake of brevity, we do not report per-category results (Ads and Cartoon) here. Refer to Appendix F for full results.
Results. Tab. 5 highlights a significant trend: models struggle to distinguish negatives within the image (local), but excel in identifying global negatives. A major challenge for most models was handling semantic negatives within the same image, as evidenced by the generally wide margin between models’ performance on global and local setups. Still, the global negative samples exhibited more pronounced distinctions based on their sampling scheme. Negatives sampled uniformly were distinguishable by most models with $\geq 90\%$ accuracy. In contrast, retrieval methods proved more challenging across the board, particularly for negatives retrieved using the text-to-text similarity model (textual), which increased the problem complexity for most models. Notably, OFA failed to follow zero-shot instructions for multiple-choice answering, scoring close to zero. Finally, we also present results for cropped ground-truth region images. Although cropped images are not lossless representations of the regions, all models exhibited significant improvements, indicating that the ability to infer relevant visual cues is indeed a critical challenge. Thus, we conclude that models struggle to infer which visual cues support the argument.
Figure 5: Failure cases of LLaVA-1.5 in Identification of Premises. The model incorrectly reasons about relevant objects, relying instead on common words.
Image $\Delta$ VP $\Delta$ CP $\Delta$ Tree
LLaVA-1.5 3.48 $\uparrow$ 13.29 (5.42) $\uparrow$ 8.57 (4.75) $\uparrow$ 4.72 (4.34)
LLaVA-Next 15.04 $\uparrow$ 11.28 (1.21) $\uparrow$ 6.72 (2.78) $\uparrow$ 4.14 (4.02)
Table 7: Mean of incremental improvements in BERTScore with each additional input across four different prompts in Deduction of Conclusion. Standard deviations are shown in parentheses.
5.3 Deduction of Conclusion
Deduction of Conclusion evaluates the comprehensive ability to deduce the conclusion of an argument. Given a subset of inputs among the image $I$, the visual premises VP, the commonsense premises CP, and the reasoning tree $T$, the objective is to generate the conclusion C of an argument.
Metrics and Models. As discussed earlier in § 4.3, we use BERTScore as the primary metric. We supplement this with three additional static metrics (Bleu-4, ROUGE-L, CIDEr) in Appendix F. The models tested in this task include all the multimodal LLMs used in the previous experiment and text-only LLMs (LLaMa-3-Instruct AI@Meta (2024), Mistral-Instruct Jiang et al. (2023), and Zephyr Tunstall et al. (2023)). All LLMs considered here are the $7\sim 8$b sized variants. The LLMs do not take the image as an input.
Results. Table 6 shows the results for this task. As expected from previous tasks, most models experience the highest gain from the additional information provided by the ground-truth set of visual premises. This supports our hypothesis that selective attention to visual premises is a bottleneck in understanding visual arguments in current models. Also, both multimodal and text-only models benefited from commonsense premises and reasoning trees in most setups, indicating that models cannot yet perfectly understand visual arguments in a text-only format and benefit from explicit reasoning process information. We note that OFA struggled to follow the instruction format, leading to sub-zero scores. Although rare, BERTScore, based on cosine similarity, can yield negative values. We also clarify that the multimodality of the deduction of conclusion task resides in the visual premises, making it solvable by text-only models given them.
5.4 Diagnostics
Prompt Robustness. To ensure the robustness of our empirical results, we differentiated the prompts provided to the models. As shown in Tab. 7, the trend of gains remained stable across four different prompts, confirming the validity of our tests. For detailed prompts, refer to Appendix M, and for results in other tasks, see Appendix E.
Error Analysis. Fig. 5 provides qualitative examples of failure cases. We present straightforward instances to clearly explain the errors. In these cases, the models fail to reason about the relevant object, which is the subject of the given intermediate conclusion, and instead rely on common words, leading to incorrect inference results.
6 Conclusion
We introduce VisArgs, a curated and annotated benchmark for visual argument understanding. Using our benchmark, we affirm a compelling hypothesis: selective vision is a critical bottleneck for visual reasoning in current machines. We aim for our benchmark to serve as a resource for advancing multimodal intelligence beyond passive captioning. Future work includes:
  1. 1.
    Conditional Saliency Analysis: It is demonstrated that the saliency required for visual arguments differs from that needed for passive captioning. Can the varying saliency requirements across different tasks be analyzed?
  2. 2.
    Extending Modalities: In speech recognition, non-conditional selective attention is known as the cocktail party effect. Would conditional selective attention be necessary in modalities other than vision as well?
7 Limitations
VisArgs, which is built on advertisements and cartoons from web sources, does not encompass all forms of visual arguments. Visual arguments also include various forms of media including mathematical diagrams Inglis and Mejía-Ramos (2009) and videos, such as films Alcolea-Banegas (2009). Consequently, the findings of this study do not represent all forms of visual arguments.
Additionally, the annotations for VisArgs are created by two NLP researchers with similar cultural backgrounds. Although a different group of human evaluators validated these annotations, future research should consider individual variances in the interpretation of visual arguments and the reasoning processes identified by reasoning trees.
Finally, we excluded images containing written text in non-English languages when curating VisArgs, as the annotators were not familiar with other languages. This limitation may confine the cultural context covered by VisArgs, thus representing only a partial depiction of visual arguments. Since the logical relations forming a visual argument can depend on culture-specific elements, this skewed distribution of images can lead to a biased understanding of visual arguments.
Appendix A Data Annotation Details
Human Resources. To ensure a comprehensive understanding of the intricate requirements of our setup and maintain consistency across annotations, two of this paper’s authors conducted the entire annotation process. Three volunteers from the NLP research community did the human evaluation.
Annotation Interface. We used a custom-built interface for efficient and convenient image annotation. The interface is depicted in Fig. 6 and  Fig. 7. Additionally, we provide a snapshot of the human evaluation interface for Identification of Premises in Fig. 8. We will open-source this interface along with the dataset.
Appendix B Analyzing Topic Diversity
Initially, we considered using the Latent Dirichlet Allocation (LDA) Blei et al. (2003) method for data visualization, following previous literature Hessel et al. (2022). However, we found that LDA based on Bag-of-Words representations could not generate meaningful clusters or labels for conclusion topics. As a solution, we developed an adaptive semantic classification technique using multimodal large language models:
Defining Class Labels. We utilize GPT-4-O. We first sample 400 sentences each for VP and C, and then feed them to GPT with the following instructions: For VP: "Give me well-balanced 10 object type classes for these texts (e.g., eating & dining, environments & landscapes, attire). Just classes." For C: "Give me well-balanced 10 classes for these texts. Just classes." After receiving the 10 classes from the GPT, we manually refine these classes into 8 classes for both VP and C.
Labelling Data. We use a pretrained language model to classify visual premises (VP) and conclusions (C) in a zero-shot manner. We provide the following input to the LLaMA-3777meta-llama/Meta-Llama-3-8B-Instruct LLM:
Classes: {}
Your task is to classify a sentence into
the given classes.
Give me just the class.
Sentence: {}
Visualization. We use the Plotly Plotly (2015) library.
Appendix C Experiment Details
model dtype #parameter version
CLIP - 623M RN50x64
CLIP - 427M ViT-L/14
CLIP - 427M ViT-L/14@336px
SigLIP - 652M large-patch16-384
AlphaCLIP - 428M clip_l14_336_grit_20m_4xe
UNINEXT-H - 775M image_joint_vit_huge_32g
LISA - 7B xinlai/LISA-7B-v1
MM-G-DINO - 343M grounding_dino_swin-l_pretrain_all
LLaVA-1.5 FP16 7B llava-1.5-7b-hf
LLaVA-NeXT FP16 7B mistral-v0.2
Idefics2 FP16 8B chatty
OFA - 470M vqa-pretrain-large
QwenVLChat BF16 9B Qwen-VL-Chat
CogVLM BF16 17B cogvlm-chat-hf
InstructBLIP FP16 7B instructblip-vicuna-7b
Unified-IO-2 - 3B uio2-xl
Table 8: Details on the models used in our experiments.
C.1 Localization of Premises
For closed-set grounding, we utilized CLIP, SigLIP, AlphaCLIP, and OFA. We measured the alignment between regions and descriptions of visual premises using image-to-text cosine similarity scores. The input regions were provided as cropped images. A model output was considered correct (True) if the similarity between the ground-truth region and the given description was the highest among all candidates; otherwise, it was marked incorrect (False).
For open-set grounding, we employed object grounding models such as MM-GDINO, UNINEXT-H, LISA, OFA, and Unified-IO-2 to directly generate bounding box coordinates. We applied a threshold of 0.35 to the outputs, merging the selected regions into the tightest rectangle union. For LISA, we converted the output segmentation mask into bounding boxes. We then calculated the Intersection over Union (IoU) score for each bounding box. To compute the accuracy metric, we used a threshold of 0.5 for binary classification over the IoU. We calculated the local mean, which is the mean per visual premise in an image, and the mean per image.
C.2 Identification of Premises
We utilized OFA, Qwen-VL-Chat, CogVLM, Idefics2, InstructBLIP, Unified-IO-2, LLaVA-1.5, LLaVA-Next, and GPT-4-O for our experiments. We created multiple-choice questions with three possible answers: one correct answer and two incorrect answers. Five conditions were set for sampling the negatives for incorrect answers:
  • Random Sampling: This global sampler selects samples uniformly without duplication.
  • Visual Sampling: This global sampler chooses the top 2 premise descriptions most similar to the image, using CLIP to score the cosine similarity between the image and text. We set the CLIP similarity threshold to 0.24 to ensure negative premises do not accurately describe the image.
  • Textual Sampling: This sampler selects the top 2 premise descriptions most similar to the ground truth premise, using ColBERT to score the cosine similarity between texts. We set the ColBERT similarity threshold to 25 to prevent negative premises from accurately describing the image.
  • Mixed Sampling: This approach combines visual and textual sampling, visually selecting from the top 10 textual retrieval results.
To ensure a fair comparison across various negative sampling methods, we use only intermediate conclusions that have three or more related visual premises. This results in 1,775 visual premises for the advertisement category and 1,774 for the cartoon category, totaling 3,549 visual premises, which is 62.34% of the overall visual premises.
Human Evaluation. We randomly selected 100 images from each data category and had human annotators perform the same tests as the machines across all negative set setups. The results demonstrated that humans achieved nearly perfect accuracy in this task, as shown in Tab. 12.
C.3 Deduction of Conclusion
We conducted experiments on both Multi-Modal Large Language Models (MLLM) and Large Language Models (LLMs). The MLLMs used in our experiments include LLaVA-1.5, LLaVA-NeXT, Idefics2, OFA, InstructBLIP, Qwen-VL-Chat, CogVLM, and Unified-IO-2. The LLMs include LLaMA-3, Mistral, and Zephyr.
Prompting. Before conducting the experiments, we established a set of instructions to be applied to all models to elicit appropriate responses. During this process, we encountered several issues with prompt engineering, such as model refusal to address controversial or unsafe questions, the inclusion of unnecessary tokens, multiple sentences, and the positioning of image tokens. Ultimately, we decided on the following prompt: "<image> <information> Your task is to answer what the image wants to convey. You should respond in only one sentence without any unnecessary prefixes. ANSWER:"
C.4 Resource & Hyperparameters
Computation. We utilized RTX-4090 and A6000 GPUs for our experiments. All models, except for CogVLM, were implemented using RTX-4090 GPUs. Due to the size of its model weights, CogVLM was implemented on an A6000 GPU. Each model required up to 8 RTX-4090 GPU-hours per task. In total, conducting all tasks demanded 200 RTX-4090 GPU-hours.
Hyperparameters. Our experiments are deterministic, given the pretrained model weights, the greedy decoding scheme, and the instruction prompts. We explore prompt diversification in § 5.4 and Appendix E.
C.5 Model Details
We specify all exact model identifiers and sizes in Tab. 8.
Appendix D Comparison of Metrics for Deduction of Conclusion
Here, we describe details for human evaluation of goodness per each metric illustrated in Tab. 2.
Human Evaluation. We sampled 200 target images and collected responses from three models: LLaVA-Next, Qwen-VL-Chat, and GPT-4o. Human annotators then determined whether each model’s conclusion was semantically similar to the reference conclusion.
Metrics. To evaluate accuracy, precision, recall, and F1-score, we first converted each metric into binary decisions using derived thresholds. We established these thresholds by training a logistic regression model on 100 pairs of metric scores and human decisions. Subsequently, we inferred binary decision labels on the remaining 100 pairs. The results are presented in Tab. 9. Additionally, the correlation between the metrics and human decisions is reported using Pearson’s coefficient Cohen et al. (2009).
Accuracy Precision Recall F1-score Pearson Corr. ($\rho$)
BLEU-4 0.67 0.44 0.67 0.53 0.18
ROUGE 0.75 0.76 0.75 0.72 0.35
CIDEr 0.72 0.70 0.72 0.70 0.26
GPTEval 0.75 0.83 0.75 0.76 0.53
BERTScore 0.94 0.94 0.93 0.93 0.59
Table 9: Comparison of metrics with human decision on Deduction of Conclusion
Appendix E Prompt Robustness in Identification of Premises
Global Local
Prompt Random Visual Textual Mixed Semantic
LLaVA-NeXT Original 97.10 96.14 80.53 84.70 77.51
InstructBLIP 90.65 84.53 71.54 74.75 58.21
LLaVA-NeXT Paraphrase 1 97.24 96.59 80.98 85.63 77.60
InstructBLIP 90.93 84.73 72.78 74.98 59.68
LLaVA-NeXT Paraphrase 2 97.60 96.25 81.46 86.00 76.67
InstructBLIP 93.32 89.83 79.01 81.71 64.50
Table 10: Assessment of prompt robustness with different paraphrases in Identification of Premises. Accuracy is measured as a percentage.
[…]
Ads Cartoon All Ads Cartoon All
UNINEXT-H 34.50 44.33 38.75 31.67 40.71 35.58
LISA 40.05 49.17 44.25 40.52 50.01 44.62
Unified-IO-2 45.81 52.29 48.61 44.66 50.43 47.15
OFA 49.10 51.49 50.14 49.06 49.22 49.13
MM-G-Dino 52.70 58.06 55.02 52.39 58.37 54.98
Table 11: Open-set grounding results in localization of premises.
Extending the robustness study in Tab. 7, we conducted a similar prompt diversification experiment for the task of Identification of Premises. By paraphrasing the original prompt as described in Appendix L, we performed the same evaluation. The results, presented in Tab. 10, demonstrate that our experimental outcomes remain stable for Identification of Premises across different prompt paraphrases.
Appendix F Full Results
This section presents the comprehensive versions of the results summarized in the main paper. Tab. 11 displays the open-set grounding results for Localization of Premises, while Tab. 12 provides the results for Identification of Premises. The results for the task of Deduction of Conclusion are detailed by category: advertisements are shown in Tab. 13, cartoons in Tab. 14, and the average across both categories in Tab. 15.
Appendix G Qualitative Samples on Open-Set Grounding
To identify the cause of low performance in the open-set evaluation of the Localization of Premises task, we examine qualitative samples shown in Fig. 9. Traditional object detection models are typically trained on single object labels, whereas our semantic region labels may encompass multiple objects with similar meanings. Consequently, although the models may detect the correct target, the intersection over union (IoU) scores are lower, resulting in reduced accuracy.
Appendix H Qualitative Samples on Deduction of Conclusion
Inference results of different models with varying inputs are shown in Fig. 10 and Fig. 11. The outputs of the models display discrepancies; for instance, CogVLM exhibits weak conditioning on additional inputs, producing similar outputs despite the incremental increase in information provided through different inputs.
Appendix I Credits
We do not claim any rights to the images included in our dataset. Therefore, we provide only the URLs to the corresponding images instead of distributing the raw files. For usage outside of an academic context, please contact the copyright holders directly.
Appendix J Prompts for Annotation
  • Annotation for Visual Premises
    Your task is to identify visual premises from the image. These are visual cues that support or illustrate the conclusion, enhancing the overall understanding and clarity of the image.
    Example
    Visual Premises (VP):
    1. The image depicts a maze with entry point and exit.
    2. At the entry point of a maze labeled "Start," there is a cigarette.
    3. The exit of the maze is labeled "Lung Cancer."
    4. Theres a text saying, "Or you can start here," with an arrow pointing to another text that reads, "Make the right choice. DONT SMOKE."
  • Annotation for Constructing Arguments
    Visual Premises (VP):
    1. VP1
    2. VP2
    3. VP3
    Given the visual premises of the image, your task is to generate the necessary commonsense premises and conclusion of the image. The conclusion should be one simple sentence. Then show the reasoning steps to reach the conclusion. The reasoning steps should include all visual premises and commonsense premises. You can refer to the following example.
    Example
    Visual Premises (VP):
    1. The image depicts a maze with entry point and exit.
    2. At the entry point of a maze labeled "Start," there is a cigarette.
    3. The exit of the maze is labeled "Lung Cancer."
    4. Theres a text saying, "Or you can start here," with an arrow pointing to another text that reads, "Make the right choice. DONT SMOKE."
    Commonsense Premises (CP):
    1. Mazes are often used to represent complex journeys or paths one must navigate.
    2. Cigarettes are known to be harmful to health and a major cause of lung cancer.
    3. The phrase "Make the right choice" implies that there is a decision to be made that can impact ones health.
    4. Public health messages often use strong visuals to convey the importance of making healthy choices.
    Conclusion (C):
    The image is a public health message that illustrates the dangerous path from smoking to lung cancer while encouraging individuals to choose not to smoke for their health.
    Reasoning Steps:
    (VP1, CP1 -> IC1): The maze represents the difficult and potentially harmful journey.
    (VP2, CP2 -> IC2): The presence of a cigarette at the mazes entry point indicates the start of this hazardous journey.
    (VP3, CP2 -> IC3): Labeling the mazes exit as "Lung Cancer" directly links smoking to this deadly disease.
    (VP4, CP3, CP4 -> IC4): The additional text offers an alternative choice to avoid smoking, emphasizing the importance of preventive health measures.
    (IC1, IC2, IC3, IC4 -> C): The image is a public health message that warns about the risks of smoking and encourages making the right choice for ones health.
    Answer
Appendix K Prompts for Evaluation
  • GPTEval
    Task Description: You will be given a ground truth sentence that describes an image and a model-generated sentence. Your task is to evaluate the semantic similarity between the model-generated sentence and the ground truth sentence. You dont need to give me any description. Just score should be answered.
    Evaluation Criteria: T/F. False means the sentences are completely different. True means they mean exactly the same thing.
    Ground Truth: {}
    Generated: {}
Appendix L Prompts for Identification of Retrieval
  • Original Prompt
    <image>
    The following are multiple choice questions (with answers) about image understanding.
    When given an image, a conclusion, and several visual cue options, you need to identify the visual cue that best relates to the conclusion. To do this effectively, carefully analyze how each visual cue connects to the key elements of the conclusion. Select the visual cue that most directly supports or illustrates the conclusion, ensuring that it enhances the overall understanding and clarity of the message. Answer A), B), or C) with no additional explanation. Conclusion: {conclusion}
    {vp_options}
    ANSWER:
  • Paraphrase 1
    <image>
    The following are multiple choice questions (with answers) about image understanding.
    When given an image, a conclusion, and several visual cue options, identify the visual cue that best relates to the conclusion. Select the visual cue that most directly supports or illustrates the conclusion, ensuring that it enhances the overall understanding and clarity of the message. To do this effectively, carefully analyze how each visual cue connects to the key elements of the conclusion. Answer A), B), or C) with no additional explanation. Conclusion: {conclusion}
    {vp_options}
    ANSWER:
  • Paraphrase 2
    <image>
    The following are multiple choice questions (with answers) about image understanding.
    Given an image, what is the visual cue most related to the given conclusion? Answer A), B), or C) with no additional explanation. Conclusion: {conclusion}
    {vp_options}
    ANSWER:
Appendix M Prompts for Deduction of Conclusion
  • Image -> C
    <image>
    Your task is to answer what the image wants to say. You should answer in only one sentence without an unnecessary prefix. ANSWER:
  • Image, VP -> C
    <image>
    "Visual Premises (VP)" are the important features presented in the images.
    Visual Premises (VP):
    1. VP1
    2. VP2
    3. VP3
    Your task is to answer what the image wants to say. You should answer in only one sentence without an unnecessary prefix. ANSWER:
  • Image, VP, CP -> C
    <image>
    "Visual Premises (VP)" are the important features presented in the images. " Commonsense Premises (CP)" are not visually depicted in the image but are commonly understood by people.
    Visual Premises (VP):
    1. VP1
    2. VP2
    3. VP3
    Commonsense Premises (CP):
    1. CP1
    2. CP2
    3. CP3
    Your task is to answer what the image wants to say. You should answer in only one sentence without an unnecessary prefix.
    ANSWER:
  • Image, VP, CP, Tree -> C
    <image>
    "Visual Premises (VP)" are the important features presented in the images. " Commonsense Premises (CP)" are not visually depicted in the image but are commonly understood by people. "Reasoning Steps" are the structure of explanation of how we came up to the "Intermediate Conclusion(IC) and "Conclusion".
    Visual Premises (VP):
    1. VP1
    2. VP2
    3. VP3
    Commonsense Premises (CP):
    1. CP1
    2. CP2
    3. CP3
    Reasoning Step:
    (VP1, CP1 -> IC1): IC1
    (VP2, CP2 -> IC2): IC2
    (VP3, CP3 -> IC3): IC3
    (IC1, IC2, IC3 -> C):
    Your task is to answer what the image wants to say. You should answer in only one sentence without an unnecessary prefix. ANSWER:
  • Prompt Style 1
    <image>
    "Visual Premises (VP)" are the important features presented in the images. " Commonsense Premises (CP)" are not visually depicted in the image but are commonly understood by people. "Reasoning Steps" are the structure of explanation of how we came up to the "Intermediate Conclusion(IC) and "Conclusion".
    Visual Premises (VP):
    Commonsense Premises (CP):
    Reasoning Step:
    Answer in one sentence what the image wants to convey. ANSWER:
  • Prompt Style 2
    <image>
    "Visual Premises (VP)" are the key visual elements in the image. "Commonsense Premises (CP)" are elements based on general common sense. "Reasoning Steps" are the process of reaching the "Intermediate Conclusion (IC)" and "Conclusion".
    Visual Premises (VP):
    Commonsense Premises (CP):
    Reasoning Step:
    Write the message of the image in one sentence. You should answer in only one sentence without an unnecessary prefix. RESPONSE:
  • Prompt Style 3
    <image>
    "Visual Premises (VP)" represent the important features observed in the image. "Commonsense Premises (CP)" are things not visually depicted but generally understood. "Reasoning Steps" are the explanation process leading to the "Intermediate Conclusion (IC)" and "Conclusion".
    Visual Premises (VP):
    Commonsense Premises (CP):
    Reasoning Step:
    Write the main message of the image in one sentence. RESPONSE:
  • Prompt Style 4
    <image>
    "Visual Premises (VP)" are the key features observed in the image. "Commonsense Premises (CP)" are not visually depicted but can be understood through general knowledge. "Reasoning Steps" are the logical explanation process leading to the "Intermediate Conclusion (IC)" and "Conclusion".
    Visual Premises (VP):
    Commonsense Premises (CP):
    Reasoning Step:
    Write the meaning the image wants to convey in one sentence. RESPONSE:
Global Local
Random Visual Textual Mixed Semantic
Ads Cartoon All Ads Cartoon All Ads Cartoon All Ads Cartoon All Ads Cartoon All
Random 33.33 33.33 33.33 33.33 33.33 33.33 33.33 33.33 33.33 33.33 33.33 33.33 33.33 33.33 33.33
Human 100.00 100.00 100.00 100.00 98.00 99.00 96.00 92.00 94.00 100.00 100.00 100.00 98.00 98.00 98.00
OFA 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Qwen-VL-Chat 88.90 83.21 86.05 88.67 82.87 85.77 73.73 67.61 70.67 77.00 74.14 49.73 53.21 46.25 75.57
CogVLM 97.58 97.35 97.46 96.45 96.34 96.39 88.78 87.21 88.00 91.66 92.79 92.22 69.28 61.35 65.31
Idefics2 98.59 98.76 98.68 97.91 97.75 97.83 93.18 90.42 91.80 95.15 94.99 95.07 77.40 72.62 75.01
InstructBLIP 82.41 85.13 83.77 78.07 80.39 79.23 68.55 65.35 66.95 71.87 70.87 71.37 66.91 56.90 61.90
Unified-IO-2 98.31 98.54 98.42 97.29 96.68 96.99 88.78 84.96 86.87 92.28 93.35 92.81 34.67 34.82 34.74
LLaVA-1.5 98.82 98.48 98.65 98.08 97.75 97.91 84.44 83.04 83.74 89.23 90.48 89.86 73.34 61.52 67.43
LLaVA-NeXT 97.35 97.97 97.66 96.05 96.34 96.20 81.17 80.62 80.90 84.33 87.38 85.86 82.69 74.37 78.53
GPT-4-O - - - - - - - - - - - - 75.22 82.56 79.50
Table 12: Results on Identification of Premises.
Inputs Automatic Semantic
I VP CP RS BLEU-4 ROUGE CIDEr BERT
LLaMA3 7.07 28.41 33.08 43.00
8.65 ($\uparrow$ 1.58) 31.44 ($\uparrow$ 3.03) 40.87 ($\uparrow$ 7.79) 59.58 ($\uparrow$ 16.58)
8.34 ($\downarrow$ 0.31) 31.18 ($\downarrow$ 0.26) 41.94 ($\uparrow$ 1.07) 56.70 ($\downarrow$ 2.88)
Mistral 2.95 19.84 23.28 24.86
4.95 ($\uparrow$ 2.00) 25.13 ($\uparrow$ 5.29) 33.90 ($\uparrow$ 10.62) 39.92 ($\uparrow$ 16.05)
6.15 ($\uparrow$ 1.20) 27.06 ($\uparrow$ 1.93) 38.34 ($\uparrow$ 4.43) 49.54 ($\uparrow$ 9.62)
Zephyr 2.78 16.35 24.06 16.15
3.25 ($\uparrow$ 0.47) 17.44 ($\uparrow$ 1.08) 30.41 ($\uparrow$ 6.35) 31.38 ($\uparrow$ 6.52)
5.20 ($\uparrow$ 1.94) 22.70 ($\uparrow$ 5.26) 36.29 ($\uparrow$ 5.88) 45.23 ($\uparrow$ 13.85)
OFA 0.00 0.13 0.01 -41.26
0.00 (-) 5.24 ($\uparrow$ 5.10) 0.47 ($\uparrow$ 0.47) -22.52 ($\uparrow$ 18.75)
0.00 (-) 5.79 ($\uparrow$ 0.55) 0.37 ($\downarrow$ 0.10) -15.87 ($\uparrow$ 6.65)
0.00 (-) 6.53 ($\uparrow$ 0.75) 0.70 ($\uparrow$ 0.33) -12.51 ($\uparrow$ 3.36)
QwenVLChat 0.72 13.12 8.41 14.32
4.02 ($\uparrow$ 3.30) 24.73 ($\uparrow$ 11.61) 30.58 ($\uparrow$ 22.17) 28.74 ($\uparrow$ 14.41)
4.85 ($\uparrow$ 0.83) 26.67 ($\uparrow$ 1.94) 35.30 ($\uparrow$ 4.72) 34.05 ($\uparrow$ 5.31)
4.89 ($\uparrow$ 0.03) 26.89 ($\uparrow$ 0.23) 38.30 ($\uparrow$ 3.00) 35.11 ($\uparrow$ 1.06)
CogVLM 4.96 24.40 25.56 27.38
6.06 ($\uparrow$ 1.09) 27.37 ($\uparrow$ 2.97) 39.19 ($\uparrow$ 13.63) 33.24 ($\uparrow$ 5.87)
7.18 ($\uparrow$ 1.13) 29.25 ($\uparrow$ 1.88) 47.21 ($\uparrow$ 8.02) 36.41 ($\uparrow$ 3.17)
7.68 ($\uparrow$ 0.50) 30.03 ($\uparrow$ 0.78) 51.44 ($\uparrow$ 4.23) 37.71 ($\uparrow$ 1.29)
Idefics2 4.00 21.97 18.56 21.27
4.53 ($\uparrow$ 0.53) 24.13 ($\uparrow$ 2.17) 28.79 ($\uparrow$ 10.24) 27.39 ($\uparrow$ 6.12)
5.56 ($\uparrow$ 1.03) 25.17 ($\uparrow$ 1.04) 38.21 ($\uparrow$ 9.42) 33.22 ($\uparrow$ 5.83)
7.48 ($\uparrow$ 1.92) 27.73 ($\uparrow$ 2.56) 53.22 ($\uparrow$ 15.01) 38.40 ($\uparrow$ 5.17)
InstructBLIP 0.00 4.22 1.01 -15.92
3.27 ($\uparrow$ 3.26) 18.94 ($\uparrow$ 16.26) 22.16 ($\uparrow$ 21.15) 23.15 ($\uparrow$ 39.07)
6.00 ($\uparrow$ 2.73) 26.91 ($\uparrow$ 7.32) 44.23 ($\uparrow$ 22.07) 35.20 ($\uparrow$ 12.05)
6.27 ($\uparrow$ 0.27) 28.29 ($\uparrow$ 0.37) 45.53 ($\uparrow$ 1.29) 35.52 ($\uparrow$ 0.32)
Unified-io 2 0.07 10.02 0.89 -8.49
0.65 ($\uparrow$ 0.58) 14.81 ($\uparrow$ 4.79) 5.18 ($\uparrow$ 4.29) -0.04 ($\uparrow$ 8.45)
0.82 ($\uparrow$ 0.17) 15.27 ($\uparrow$ 0.46) 7.73 ($\uparrow$ 2.55) 5.63 ($\uparrow$ 5.68)
0.93 ($\uparrow$ 0.11) 16.10 ($\uparrow$ 0.83) 10.12 ($\uparrow$ 2.39) 8.43 ($\uparrow$ 2.79)
LLaVA 1.38 14.93 3.73 3.25
3.92 ($\uparrow$ 2.54) 22.93 ($\uparrow$ 8.01) 21.49 ($\uparrow$ 17.76) 22.21 ($\uparrow$ 18.96)
5.49 ($\uparrow$ 1.58) 26.40 ($\uparrow$ 3.47) 39.29 ($\uparrow$ 17.80) 32.48 ($\uparrow$ 10.28)
5.68 ($\uparrow$ 0.19) 26.46 ($\uparrow$ 0.06) 42.65 ($\uparrow$ 3.36) 34.22 ($\uparrow$ 1.74)
LLaVA-NeXT 3.62 21.08 15.23 18.05
6.78 ($\uparrow$ 3.16) 28.31 ($\uparrow$ 7.23) 42.17 ($\uparrow$ 26.94) 32.98 ($\uparrow$ 14.94)
7.51 ($\uparrow$ 0.73) 30.03 ($\uparrow$ 1.72) 50.54 ($\uparrow$ 8.37) 37.93 ($\uparrow$ 4.95)
8.51 ($\uparrow$ 1.00) 31.19 ($\uparrow$ 1.15) 61.70 ($\uparrow$ 11.16) 40.96 ($\uparrow$ 3.02)
GPT-4-O 2.38 17.20 23.08 25.96
5.44 ($\uparrow$ 3.07) 24.47 ($\uparrow$ 7.27) 46.05 ($\uparrow$ 22.98) 36.09 ($\uparrow$ 10.13)
6.82 ($\uparrow$ 1.38) 26.36 ($\uparrow$ 1.88) 61.20 ($\uparrow$ 15.15) 41.08 ($\uparrow$ 4.99)
Table 13: Results on Deduction of Conclusion in the Advertisement category.
Inputs Automatic Semantic
I VP CP RS BLEU-4 ROUGE CIDEr BERT
LLaMA3 5.51 27.67 31.28 26.46
6.65 ($\uparrow$ 1.13) 29.95 ($\uparrow$ 2.28) 48.01 ($\uparrow$ 16.73) 33.71 ($\uparrow$ 7.25)
8.81 ($\uparrow$ 2.17) 32.17 ($\uparrow$ 2.22) 61.74 ($\uparrow$ 13.73) 39.20 ($\uparrow$ 5.49)
Mistral 1.87 16.29 11.24 13.23
4.01 ($\uparrow$ 2.14) 22.24 ($\uparrow$ 5.95) 28.15 ($\uparrow$ 16.91) 25.22 ($\uparrow$ 11.99)
6.45 ($\uparrow$ 2.44) 27.82 ($\uparrow$ 5.59) 45.00 ($\uparrow$ 16.85) 34.39 ($\uparrow$ 9.17)
Zephyr 1.70 13.88 11.54 16.15
3.28 ($\uparrow$ 1.58) 17.51 ($\uparrow$ 3.63) 24.92 ($\uparrow$ 13.38) 26.40 ($\uparrow$ 10.25)
6.91 ($\uparrow$ 3.63) 27.20 ($\uparrow$ 9.69) 46.63 ($\uparrow$ 21.71) 36.70 ($\uparrow$ 10.29)
OFA 0.00 0.45 0.01 -41.35
0.00 (-) 5.30 ($\uparrow$ 4.84) 0.27 ($\uparrow$ 0.26) -27.23 ($\uparrow$ 14.12)
0.00 (-) 8.15 ($\uparrow$ 2.85) 0.26 ($\downarrow$ 0.01) -17.30 ($\uparrow$ 9.93)
0.00 (-) 8.45 ($\uparrow$ 0.30) 0.58 ($\uparrow$ 0.32) -15.74 ($\uparrow$ 1.56)
QwenVLChat 0.49 13.98 4.23 10.79
3.18 ($\uparrow$ 2.69) 23.80 ($\uparrow$ 9.81) 18.02 ($\uparrow$ 13.79) 17.18 ($\uparrow$ 6.39)
4.23 ($\uparrow$ 1.05) 26.40 ($\uparrow$ 2.60) 26.76 ($\uparrow$ 8.74) 25.03 ($\uparrow$ 7.85)
4.79 ($\uparrow$ 0.56) 27.87 ($\uparrow$ 1.47) 32.10 ($\uparrow$ 5.34) 29.49 ($\uparrow$ 4.46)
CogVLM 4.89 27.48 24.56 23.51
5.49 ($\uparrow$ 0.60) 28.50 ($\uparrow$ 1.02) 32.11 ($\uparrow$ 7.54) 27.24 ($\uparrow$ 3.73)
6.43 ($\uparrow$ 0.94) 29.64 ($\uparrow$ 1.14) 38.02 ($\uparrow$ 5.91) 29.88 ($\uparrow$ 2.64)
7.89 ($\uparrow$ 1.46) 31.72 ($\uparrow$ 2.08) 53.05 ($\uparrow$ 15.04) 34.45 ($\uparrow$ 4.56)
Idefics2 3.50 21.74 16.07 16.36
3.61 ($\uparrow$ 0.78) 23.47 ($\uparrow$ 2.04) 20.02 ($\uparrow$ 7.20) 16.70 ($\uparrow$ 6.78)
5.42 ($\uparrow$ 1.81) 26.58 ($\uparrow$ 3.11) 29.96 ($\uparrow$ 9.94) 24.50 ($\uparrow$ 7.81)
8.28 ($\uparrow$ 2.86) 31.01 ($\uparrow$ 4.43) 54.64 ($\uparrow$ 24.68) 34.27 ($\uparrow$ 9.76)
InstructBLIP 0.00 3.24 0.40 -21.55
2.53 ($\uparrow$ 2.53) 18.94 ($\uparrow$ 15.14) 16.35 ($\uparrow$ 15.60) 16.62 ($\uparrow$ 34.98)
5.33 ($\uparrow$ 2.80) 26.91 ($\uparrow$ 7.98) 36.87 ($\uparrow$ 20.52) 28.92 ($\uparrow$ 12.30)
6.18 ($\uparrow$ 0.85) 28.29 ($\uparrow$ 1.38) 42.53 ($\uparrow$ 5.65) 32.17 ($\uparrow$ 3.25)
Unified-io 2 0.00 9.07 0.40 -11.68
0.56 ($\uparrow$ 0.56) 11.32 ($\uparrow$ 2.25) 2.60 ($\uparrow$ 2.20) -7.80 ($\uparrow$ 3.88)
0.62 ($\uparrow$ 0.06) 13.79 ($\uparrow$ 2.48) 5.40 ($\uparrow$ 2.80) 2.39 ($\uparrow$ 10.19)
1.33 ($\uparrow$ 0.71) 16.50 ($\uparrow$ 2.70) 12.63 ($\uparrow$ 7.24) 7.47 ($\uparrow$ 5.08)
LLaVA 1.46 17.84 3.55 0.92
3.73 ($\uparrow$ 2.27) 23.22 ($\uparrow$ 5.38) 12.56 ($\uparrow$ 9.01) 14.92 ($\uparrow$ 14.01)
5.61 ($\uparrow$ 1.87) 27.63 ($\uparrow$ 4.41) 28.99 ($\uparrow$ 16.44) 25.89 ($\uparrow$ 10.97)
7.87 ($\uparrow$ 2.26) 30.46 ($\uparrow$ 2.84) 47.25 ($\uparrow$ 18.25) 33.11 ($\uparrow$ 7.22)
LLaVA-NeXT 2.71 22.32 11.26 11.26
5.78 ($\uparrow$ 3.07) 27.79 ($\uparrow$ 5.48) 28.12 ($\uparrow$ 16.86) 22.46 ($\uparrow$ 11.20)
7.31 ($\uparrow$ 1.52) 30.44 ($\uparrow$ 2.65) 43.84 ($\uparrow$ 15.73) 29.57 ($\uparrow$ 7.11)
9.16 ($\uparrow$ 1.85) 33.09 ($\uparrow$ 2.65) 61.44 ($\uparrow$ 17.60) 35.88 ($\uparrow$ 6.32)
GPT-4-O 4.07 23.37 23.15 24.96
6.40 ($\uparrow$ 2.34) 27.39 ($\uparrow$ 4.02) 40.64 ($\uparrow$ 17.50) 32.05 ($\uparrow$ 7.09)
8.69 ($\uparrow$ 2.29) 31.32 ($\uparrow$ 3.93) 63.13 ($\uparrow$ 22.48) 40.78 ($\uparrow$ 8.73)
Table 14: Results on Deduction of Conclusion in the Cartoon category.
Inputs Automatic Semantic
I VP CP RS BLEU-4 ROUGE CIDEr BERT
LLaMA3 6.40 28.09 37.93 30.22
7.78 ($\uparrow$ 1.39) 30.80 ($\uparrow$ 2.71) 54.57 ($\uparrow$ 16.65) 37.77 ($\uparrow$ 7.55)
8.54 ($\uparrow$ 0.76) 31.61 ($\uparrow$ 0.82) 58.88 ($\uparrow$ 4.31) 40.75 ($\uparrow$ 2.98)
Mistral 2.48 18.30 18.41 18.93
4.54 ($\uparrow$ 2.06) 23.88 ($\uparrow$ 5.57) 34.83 ($\uparrow$ 16.42) 30.15 ($\uparrow$ 11.21)
6.28 ($\uparrow$ 1.74) 27.39 ($\uparrow$ 3.51) 47.58 ($\uparrow$ 12.75) 36.63 ($\uparrow$ 6.48)
Zephyr 2.31 15.28 19.10 20.64
3.26 ($\uparrow$ 0.95) 17.47 ($\uparrow$ 2.18) 28.59 ($\uparrow$ 9.49) 28.67 ($\uparrow$ 8.04)
5.94 ($\uparrow$ 2.67) 24.65 ($\uparrow$ 7.18) 45.84 ($\uparrow$ 17.25) 36.47 ($\uparrow$ 7.79)
OFA 0.00 0.27 0.01 -41.30
0.00 (-) 5.26 ($\uparrow$ 4.99) 0.39 ($\uparrow$ 0.38) -24.55 ($\uparrow$ 5.68)
0.00 (-) 6.81 ($\uparrow$ 1.55) 0.32 ($\downarrow$ 0.06) -16.49 ($\uparrow$ 1.73)
0.00 (-) 7.36 ($\uparrow$ 0.55) 0.65 ($\uparrow$ 0.32) -13.91 ($\uparrow$ 1.22)
QwenVLChat 0.62 13.50 6.60 12.79
3.66 ($\uparrow$ 3.04) 24.33 ($\uparrow$ 10.83) 25.15 ($\uparrow$ 18.54) 23.74 ($\uparrow$ 10.94)
4.58 ($\uparrow$ 0.92) 26.55 ($\uparrow$ 2.23) 31.61 ($\uparrow$ 6.46) 30.15 ($\uparrow$ 6.41)
4.85 ($\uparrow$ 0.26) 27.31 ($\uparrow$ 0.76) 35.62 ($\uparrow$ 4.01) 32.68 ($\uparrow$ 2.53)
Idefics2 3.50 21.74 16.07 16.36
4.13 ($\uparrow$ 0.64) 23.84 ($\uparrow$ 2.11) 25.00 ($\uparrow$ 8.92) 22.76 ($\uparrow$ 6.41)
5.50 ($\uparrow$ 1.37) 25.78 ($\uparrow$ 1.93) 34.64 ($\uparrow$ 9.64) 29.45 ($\uparrow$ 6.69)
7.82 ($\uparrow$ 2.32) 29.15 ($\uparrow$ 3.37) 53.84 ($\uparrow$ 19.20) 36.61 ($\uparrow$ 7.16)
InstructBLIP 0.00 3.80 0.75 -18.36
2.53 ($\uparrow$ 2.53) 18.94 ($\uparrow$ 15.14) 16.35 ($\uparrow$ 15.60) 16.62 ($\uparrow$ 34.98)
5.33 ($\uparrow$ 2.80) 26.91 ($\uparrow$ 7.98) 36.87 ($\uparrow$ 20.52) 28.92 ($\uparrow$ 12.30)
6.18 ($\uparrow$ 0.85) 28.29 ($\uparrow$ 1.38) 42.53 ($\uparrow$ 5.65) 32.17 ($\uparrow$ 3.25)
CogVLM 4.93 25.73 25.13 25.53
5.81 ($\uparrow$ 0.88) 27.86 ($\uparrow$ 2.13) 36.13 ($\uparrow$ 11.00) 30.65 ($\uparrow$ 4.94)
6.86 ($\uparrow$ 1.04) 29.42 ($\uparrow$ 1.56) 43.23 ($\uparrow$ 7.10) 33.59 ($\uparrow$ 2.94)
7.77 ($\uparrow$ 0.92) 30.76 ($\uparrow$ 1.34) 52.14 ($\uparrow$ 8.90) 36.30 ($\uparrow$ 2.71)
Unified-io 2 0.04 9.61 0.68 -9.87
0.61 ($\uparrow$ 0.57) 13.30 ($\uparrow$ 3.69) 4.07 ($\uparrow$ 3.39) -3.40 ($\uparrow$ 6.47)
0.74 ($\uparrow$ 0.12) 14.63 ($\uparrow$ 1.33) 6.72 ($\uparrow$ 2.66) 4.23 ($\uparrow$ 7.63)
1.10 ($\uparrow$ 0.37) 16.27 ($\uparrow$ 1.64) 11.21 ($\uparrow$ 4.48) 8.01 ($\uparrow$ 3.78)
LLaVA 1.50 16.01 3.65 2.24
3.86 ($\uparrow$ 2.36) 22.88 ($\uparrow$ 6.87) 18.69 ($\uparrow$ 15.04) 19.98 ($\uparrow$ 17.74)
5.54 ($\uparrow$ 1.69) 26.93 ($\uparrow$ 4.05) 34.84 ($\uparrow$ 16.14) 29.63 ($\uparrow$ 9.66)
6.63 ($\uparrow$ 1.09) 28.19 ($\uparrow$ 1.26) 44.64 ($\uparrow$ 9.80) 33.74 ($\uparrow$ 4.11)
LLaVA-NeXT 3.23 21.62 13.51 15.11
6.35 ($\uparrow$ 3.12) 28.09 ($\uparrow$ 6.47) 36.09 ($\uparrow$ 22.58) 28.43 ($\uparrow$ 16.75)
7.42 ($\uparrow$ 1.07) 30.21 ($\uparrow$ 2.12) 47.64 ($\uparrow$ 11.55) 34.31 ($\uparrow$ 8.07)
8.46 ($\uparrow$ 1.04) 31.69 ($\uparrow$ 1.49) 61.14 ($\uparrow$ 13.50) 39.50 ($\uparrow$ 2.58)
GPT-4-O 3.11 19.87 23.11 24.50
5.86 ($\uparrow$ 2.75) 25.74 ($\uparrow$ 5.87) 43.71 ($\uparrow$ 20.61) 31.89 ($\uparrow$ 7.39)
7.63 ($\uparrow$ 1.77) 28.51 ($\uparrow$ 2.77) 62.03 ($\uparrow$ 18.32) 38.42 ($\uparrow$ 6.53)
Table 15: Results for the Deduction of Conclusion averaged across the two categories.
Figure 6: Human annotation interface for collecting textual annotations.
Figure 7: Human annotation interface for collecting bounding boxes of visual premises.
Figure 8: Human evaluation interface for Identification of Premises.
Figure 9: Qualitative samples of open-set grounding results.
Figure 10: Qualitative samples of Deduction of Conclusion, with inference results from LLaVA-NeXT and InstructBLIP models.
Figure 11: Qualitative samples of Deduction of Conclusion, with inference results from CogVLM and Qwen-VL-Chat models.

### v3

@@
 While visual arguments are readily appreciated by human audiences, we ask: are today’s AI capable of similar understanding?
-We collect and release VisArgs, an annotated corpus designed to make explicit the (usually implicit) structures underlying visual arguments.
-VisArgs includes 1,611 images accompanied by three types of textual annotations: 5,112 visual premises (with region annotations), 5,574 commonsense premises, and reasoning trees connecting them to a broader argument.
-We propose three tasks over VisArgs to probe machine capacity for visual argument understanding: localization of premises, identification of premises, and deduction of conclusions.
-Experiments demonstrate that 1) machines cannot fully identify the relevant visual cues.
-The top-performing model, GPT-4-O, achieved an accuracy of only 78.5%, whereas humans reached 98.0%.
-All models showed a performance drop, with an average decrease in accuracy of 19.5%, when the comparison set was changed from objects outside the image to irrelevant objects within the image.
-Furthermore, 2) this limitation is the greatest factor impacting their performance in understanding visual arguments.
-Most models improved the most when given relevant visual premises as additional inputs, compared to other inputs, for deducing the conclusion of the visual argument.
-Our code and data are publicly released at https://github.com/JiwanChung/VisArgs.
+We present VisArgs  [^1] Data: https://huggingface.co/datasets/jiwan-chung/visargs, a dataset of 1,611 images annotated with 5,112 visual premises (with regions), 5,574 commonsense premises, and reasoning trees connecting them into structured arguments.
+We propose three tasks for evaluating visual argument understanding: premise localization, premise identification, and conclusion deduction.
+Experiments [^2] Code: https://github.com/JiwanChung/VisArgs show that 1) machines struggle to capture visual cues: GPT-4-O achieved 78.5% accuracy, while humans reached 98.0%.
+Models also performed 19.5% worse when distinguishing between irrelevant objects within the image compared to external objects.
+2) Providing relevant visual premises improved model performance significantly.
 Selective Vision is the Challenge for Visual Reasoning:
@@
 Examples of visual arguments are prevalent in advertisements Kjeldsen (2012); Zhang et al.
+(2018); Hussain et al.
 (2017); Ye et al.
@@
 (2016); Alikhani and Stone (2018).
+Liu et al.
+Liu et al.
+(2022) also investigate arguments conveyed through images.
+However, our work focuses on visual argument, requiring that the argumentative content be primarily communicated through visual elements rather than relying solely on accompanying captions or written text.
+Furthermore, we provide explicit annotations of the argumentation structure in a tree format, facilitating detailed, hierarchical analysis of the model’s ability to comprehend visual arguments step-by-step.
 Multimodal reasoning.
@@
 Collecting Images.
-We manually collect around 1,600 images from Pinterest.222www.pinterest.com Starting with keyword-based searches (e.g.
+Our primary criterion was to select images that enable human annotators to easily and accurately interpret both the visual premise and the corresponding conclusion, thereby clarifying the argumentative structure within the image.
+Also, we ruled out samples with scene text within the images that directly describe the conclusion.
+We manually collect around 1,600 images following these criteria from Pinterest.444www.pinterest.com Starting with keyword-based searches (e.g.
 creative ads), we expanded our collection by exploring related images.
-Cartoons (which often contain visual arguments Birdsell and Groarke (1996)) were sourced from a dedicated website.333www.cartoonmovement.com We manually collected around 1,600 cartoons from various categories, including politics, education, and environment.
-For both categories, we followed previous work Schuhmann et al.
+Cartoons (which often contain visual arguments Birdsell and Groarke (1996)) were sourced from a dedicated website.555www.cartoonmovement.com We manually collected around 1,600 cartoons from various categories, including politics, education, and environment.
+We include URLs to the images to comply with licensing terms following previous work Schuhmann et al.
 (2022); Lee et al.
-(2021) by including URLs to the images to comply with licensing terms.
+(2021).
 Refer to Appendix A for details.
@@
 3.
+The human annotators could optionally incorporate new visual premises when necessary: $\sim 21\%$ of images had their set of visual premises expanded through this process.
 To facilitate this process, we break down the annotation into two steps: describing the visual premises and specifying the argument structure.
@@
 Figure 4: Variety of the topics represented in the visual premises and conclusions in VisArgs.
+Number of Samples Image Size (pixels)
+Category Total W/ Text Width Height
+Ads 914 389 877.2 969.7
+Cartoons 697 218 480.0 427.6
+Table 1: Overview of dataset statistics. W/ Text indicates the subset of images containing scene text.
 Recall Hit rate
@@
 (2023)), and LLaVA-LLaMa3 XTuner Contributors (2023) fine-tuned on a detailed captioning corpus (DOCCI Onoe et al.
-(2024))444huggingface.co/gokaygokay/llava-llama3-docci.
+(2024))666huggingface.co/gokaygokay/llava-llama3-docci.
 Tab.
@@
 Since we did not initially filter for safety, we now analyze the safety of VisArgs using standard models.
-For textual safety, we utilize the Perspective API555www.perspectiveapi.com; June 2024 version., and for visual domains, we employ LAION-Safety666www.github.com/LAION-AI/LAION-SAFETY.
+For textual safety, we utilize the Perspective API777www.perspectiveapi.com; June 2024 version., and for visual domains, we employ LAION-Safety888www.github.com/LAION-AI/LAION-SAFETY.
 The toxicity scores for textual descriptions were $0.03$ for visual premises and $0.07$ for conclusions.
@@
 We conduct a small-scale comparison study (see Tab.
-2) to verify that the model-based metric BERTScore Zhang* et al.
+3) to verify that the model-based metric BERTScore Zhang* et al.
 (2020) provides the most stable estimate, making it our primary metric.
@@
 Metrics and Models.
+We define open-set evaluation as a setting in which models are required to generate bounding box coordinates without relying on a predefined candidate list.
+As a result, models used for open-set and closed-set evaluations are architecturally distinct, since models lacking an explicit generative head, such as CLIP Radford et al.
+(2021), are not compatible with open-set evaluation due to their dependence on a candidate region list for text-to-region matching.
 For closed-set grounding, which is an N-way classification task, the goal is to match the given description with the correct bounding box.
@@
 (2022).
+The CLIP-based models are adapted as follows: for each candidate object region (specified by bounding box coordinates), the corresponding regions are cropped from the image to create region-level image representations.
+Image features for each cropped region are then extracted using a CLIP-based encoder, while text features are obtained by encoding the input description using the same model.
+Cosine similarity is calculated between the text feature and each region-level image feature.
+The region with the highest similarity score is selected as the predicted match.
 For open-set grounding, which is to locate an object in an image based on a natural language expression, we instruct the models to output bounding box coordinates and we compare them to the ground truth region.
@@
 However, the results for open-set grounding, shown in Tab.
-4, are somewhat mixed: the scores are acceptable but not uniformly high.
+5, are somewhat mixed: the scores are acceptable but not uniformly high.
 We traced this performance decline to the nature of zero-shot object detectors, which are designed to detect concrete objects and clear segments.
@@
 As shown in Tab.
-7, the trend of gains remained stable across four different prompts, confirming the validity of our tests.
+8, the trend of gains remained stable across four different prompts, confirming the validity of our tests.
 For detailed prompts, refer to Appendix M, and for results in other tasks, see Appendix E.
@@
 In these cases, the models fail to reason about the relevant object, which is the subject of the given intermediate conclusion, and instead rely on common words, leading to incorrect inference results.
+Reliance on OCR Capabilites.
+To examine how demanding VisArgs is on OCR capabilities, we employ a lightweight OCR detector Du et al.
+(2021) to detect bounding boxes of the visual premises without leveraging text annotations as input.
+Even this simplified model achieves $82.77\%$ accuracy in image-wise evaluation, where an image is considered correctly detected only if all visual premises within it are identified.
+Typical failure cases are illustrated in Fig.
+6.
+Figure 6: Left: OCR detection results.
+Right: Ground truth text instances missed by the model (highlighted in red).
+Most detection errors are attributed to Out-of-Domain cases such as calligraphy, handwritten text, or text that is too small for the model to detect, despite being distinguishable to the human eye.
 6 Conclusion
@@
 Would conditional selective attention be necessary in modalities other than vision as well?
-7 Limitations
+Limitations
 VisArgs, which is built on advertisements and cartoons from web sources, does not encompass all forms of visual arguments.
@@
 Since the logical relations forming a visual argument can depend on culture-specific elements, this skewed distribution of images can lead to a biased understanding of visual arguments.
+We encourage future research to extend this work by exploring a wider range of visual arguments and incorporating more diverse cultural and linguistic contexts.
+Acknowledgment
+This work was partly supported by an IITP grant funded by the Korean Government (MSIT) (No.
+RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University) and RS-2024-00353131) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.
+RS-2024-00354218).
 Appendix A Data Annotation Details
@@
 7 and  Fig.
-7.
+8.
 Additionally, we provide a snapshot of the human evaluation interface for Identification of Premises in Fig.
-8.
+9.
 We will open-source this interface along with the dataset.
+Inter-Annotator Agreement.
+The dataset annotation was conducted by two primary human annotators, with a third evaluator assigned to assess annotation quality and consistency.
+To measure inter-annotator agreement, 100 samples were randomly selected and re-annotated by a secondary annotator different from the original.
+Subsequently, the third evaluator independently assessed the equivalence of each annotated sample.
+Given that the annotations comprise premise sets and natural language conclusions, rather than numerical scores, traditional metrics such as Cohen’s kappa are not applicable.
+Instead, we measured inter-annotator agreement using two distinct criteria: equivalence of visual premise sets and equivalence of natural language conclusions.
+The Jaccard similarity index was employed to quantify the overlap between visual premise sets, while a binary equivalence test was used to evaluate alignment in the conclusions.
+The results, as presented in Tab.
+9, demonstrate a high degree of inter-annotator agreement.
+Qualitative analysis revealed that observed discrepancies primarily stemmed not from substantive semantic variations but from differences in how annotators segmented a single concept into one or more visual premises.
 Appendix B Analyzing Topic Diversity
@@
 We use a pretrained language model to classify visual premises (VP) and conclusions (C) in a zero-shot manner.
-We provide the following input to the LLaMA-3777meta-llama/Meta-Llama-3-8B-Instruct LLM:
+We provide the following input to the LLaMA-3999meta-llama/Meta-Llama-3-8B-Instruct LLM:
 Classes: {}
@@
 Appendix C Experiment Details
+Visual Premise Conclusion
+Human-Human 0.78 0.96
+Machine-Machine 0.51 0.88
+Table 9: Inter-annotator agreement results based on human evaluation, quantified using the Jaccard Similarity Index for set-level comparisons.
 model dtype #parameter version
@@
 The results demonstrated that humans achieved nearly perfect accuracy in this task, as shown in Tab.
-12.
+14.
 C.3 Deduction of Conclusion
@@
 The results are presented in Tab.
-9.
+11.
 Additionally, the correlation between the metrics and human decisions is reported using Pearson’s coefficient Cohen et al.
@@
 Extending the robustness study in Tab.
-7, we conducted a similar prompt diversification experiment for the task of Identification of Premises.
+8, we conducted a similar prompt diversification experiment for the task of Identification of Premises.
 By paraphrasing the original prompt as described in Appendix L, we performed the same evaluation.
 The results, presented in Tab.
-10, demonstrate that our experimental outcomes remain stable for Identification of Premises across different prompt paraphrases.
+12, demonstrate that our experimental outcomes remain stable for Identification of Premises across different prompt paraphrases.
 Appendix F Full Results
@@
 The results for the task of Deduction of Conclusion are detailed by category: advertisements are shown in Tab.
-13, cartoons in Tab.
-14, and the average across both categories in Tab.
-15.
+15, cartoons in Tab.
+16, and the average across both categories in Tab.
+17.
 Appendix G Qualitative Samples on Open-Set Grounding
 To identify the cause of low performance in the open-set evaluation of the Localization of Premises task, we examine qualitative samples shown in Fig.
-9.
+10.
 Traditional object detection models are typically trained on single object labels, whereas our semantic region labels may encompass multiple objects with similar meanings.
@@
 11 and Fig.
-11.
+12.
 The outputs of the models display discrepancies; for instance, CogVLM exhibits weak conditioning on additional inputs, producing similar outputs despite the incremental increase in information provided through different inputs.